cookiemonster:

Create a factorable polynomial with a GCF of 3x. Rewrite that polynomial in two other equivalent forms. Explain how each form was created.

Allison:

@Vocaloid @dude

cookiemonster:

please help @dude just with this one question

PandaSurvive:

This person explained it better than I could. The GCF is the greatest common factor of 2 or more numbers or polynomials. Start with the factors: x, x + 1, x - 2 Now multiply them together. x(x+1)(x−2)=(x2+x)(x−2) =x3−2x2+x2−2x =x3−x2−2x Our polynomial is: x3−x2−2x Since we created our own polynomial from factors, now we know that this polynomial factors as x3−x2−2x=x(x+1)(x−2) You can combine x with x + 1, and later x with x - 2, and have two other forms of the polynomial: x3−x2−2x=(x2+x)(x−2) x3−x2−2x=(x2−2x)(x+1)

PandaSurvive:

1 attachment
cookiemonster:

ok and you answered that in ur own words?

cookiemonster:

oh i see you copied it from the pic .. but i need u to give me a different example please i cant use that one